9. One of two supplementary angles is 30 degrees greater than twice the other. Find the degree measure of each angle.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

2A+30 + A = 180 degs
A = 50
--> 50 & 130

Answer by stanbon(75887)

One of two supplementary angles is 30 degrees greater than twice the other. Find the degree measure of each angle.
-------
1st angle: x
2nd angle: 2x+30
---
Equation:
x + 2x+30 = 180
3x = 150
x = 50 (1st angle)
2x+30 = 130 (2nd angle)
=======================
Cheers,
Stan H.
===============